# JL1-CD: A New Benchmark for Remote Sensing Change Detection and a Robust Multi-Teacher Knowledge Distillation Framework

### v3

@@ Appendix B Benchmark Methods
Appendix C Implementation Details
@@
    Training begins with a warm-up phase of 1k iterations (5 epochs for TTP), during which the learning rate (LR) is linearly increased from 1e-6 to the initial LR value, as specified in Table 6. Afterward, the LR is linearly decayed to 0 as training progresses (TTP employs a CosineAnnealing decay schedule).
@@
For the training of $\mathcal{M}_{T_{S}}$, due to the varying distributions of the change maps and the different magnitudes of $L_{CE}$ across models, a grid search is performed to determine the optimal distillation loss weight $\lambda$ for each model. More configuration details are summarized in Table 6.
Appendix D Evaluation Metrics
$\text{mIoU}=\frac{1}{2}\left(\text{IoU}_{0}+\text{IoU}_{1}\right)=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FP}+\text{FN}}+\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}}\right),$ (6)
$\text{mPrec}=\frac{1}{2}\left(\text{Prec}_{0}+\text{Prec}_{1}\right)=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FN}}+\frac{\text{TP}}{\text{TP}+\text{FP}}\right),$
$\text{mRec}=\frac{1}{2}\left(\text{Rec}_{0}+\text{Rec}_{1}\right)=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FP}}+\frac{\text{TP}}{\text{TP}+\text{FN}}\right),$
$\text{mFscore}=\frac{1}{2}\left(\text{Fscore}_{0}+\text{Fscore}_{1}\right)=\frac{1}{2}\sum_{i=0}^{1}\frac{2\times\text{Prec}_{i}\times\text{Rec}_{i}}{\text{Prec}_{i}+\text{Rec}_{i}},$
Appendix E Additional Experiments on the JL1-CD Dataset
We summarize the quantitative results of 21 models on both change and no-change regions in Table 7 to provide a comprehensive evaluation of the JL1-CD dataset as a new benchmark and to further validate the effectiveness and generality of the MTKD framework. All models exhibit improvements in IoU and Fscore for change regions after applying MTKD.
@@ Appendix E Additional Experiments on the JL1-CD Dataset
Interestingly, we observe that MTKD tends to improve recall more than precision. For instance, the recall of the change region in the MTKD-enhanced TTP model improves by 5.69%, while precision drops by 3.99%. Nevertheless, the final F1-score still increases by 3.39%. This suggests that MTKD effectively reduces the model’s miss rate, and although it may increase false alarms for some algorithms, the overall performance gains are evident.
In Figure 8, we further visualize the mIoU performance of all benchmark methods across different CAR partitions. Consistent with the findings in Section 5, MTKD yields more notable improvements in partitions with low CAR values, while its impact on high-CAR partitions is generally less pronounced. This is exactly what we need: original models already perform well on large-scale changes (with mIoU often exceeding 70% in the highest CAR partition), but struggle with small-scale changes (where mIoU frequently falls below 20% in the lowest CAR partition). MTKD makes substantial contributions to improving the detection of small-scale changes.
Figure 9 presents additional visualization results, where it can be observed that missed detections (particularly for small object changes) occur most frequently across the original algorithms. This observation aligns with the quantitative metrics reported in Table 7. Furthermore, certain algorithms (e.g., Changer-ResNet-18, CGNet, and TTP) exhibit false alarms. Both the O-P and MTKD frameworks can effectively mitigate these issues.
Whether in Table 7 or Figure 8, we also observe that some models perform worse with the O-P strategy even compared to their original versions. As previously discussed, this can be attributed to partitioning errors caused by wrong CAR estimation, which may lead to suboptimal model selection during the second-stage inference. MTKD not only improves overall test efficiency compared to O-P but also enhances model robustness. In the following section, we provide a deeper analysis of failure scenarios under the O-P strategy.
@@
Table 7: Experimental results on JL1-CD test set. To better illustrate the impact of the O-P strategy and MTKD framework on the detection performance of change and no-change classes, we report only the relative improvements in evaluation metrics compared to the original baseline model.
@@
BIT MTKD changed +2.86 +1.37 +2.09 +2.79 ChangeStar (FarSeg) MTKD changed +0.65 +0.51 +2.44 +0.82
@@
ChangeStar (UPerNet) MTKD changed +0.57 +2.31 -1.20 +0.81 ChangeFormer (MiT-b0) MTKD changed +0.28 -0.60 +0.69 +0.40
@@
ChangeFormer (MiT-b1) MTKD changed +1.68 +1.35 -0.11 +1.86 TinyCD MTKD changed +2.72 +4.13 +0.16 +2.85
@@
HANet MTKD changed +7.27 +8.96 +1.29 +8.53 Changer (MiT-b0) MTKD changed +0.80 -0.48 +2.32 +0.41
@@
Changer (MiT-b1) MTKD changed +0.41 +1.63 -1.47 +0.42 Changer (ResNet-18) MTKD changed +2.07 +4.57 -4.26 +2.57
@@
Changer (ResNeSt-50) MTKD changed +0.89 +0.36 +1.77 +1.12 LightCDNet MTKD changed +2.19 +1.70 +1.81 +2.25
@@
(a) FC-EF
(b) FC-Siam-Conc
(c) FC-Siam-Conc
(d) STANet
(e) IFN
(f) SNUNet
(g) BIT
(h) ChangeStar (FarSeg)
(i) ChangeStar (UPerNet)
(j) ChangeFormer (MiT-b0)
(k) ChangeFormer (MiT-b1)
(l) TinyCD
(m) HANet
(n) Changer (MiT-b0)
(o) Changer (MiT-b1)
(p) Changer (ResNet-18)
(q) Changer (ResNeSt-50)
(r) LightCDNet
(s) CGNet
(t) BAN
(u) TTP
(a) FC-EF
(b) STANet
(c) TTP
Appendix F Failure Case Analysis
Figure 10 presents the classification matrix, CAR distribution comparison, and scatter plot between the coarse-predicted CAR of the original models and the GT CAR under the O-P strategy for FC-EF, STANet, and TTP. For FC-EF, the scatter plot reveals a large number of points where high-GT-CAR samples are mispredicted to have almost zero CAR, which indicates a severe issue of missed detections. This observation is further corroborated by the confusion matrix, where a substantial portion of “medium” and “large” regions are incorrectly classified as “small”. As a result, the predicted CAR distribution is heavily skewed, with the number of predicted “small” samples far exceeding that of the ground truth. In contrast, STANet exhibits the opposite trend: its original model tends to generate false alarms: approximately one-quarter of “small” samples are misclassified as “medium”, and more “medium” samples are misclassified as “large” than as “small”. Such misclassification of CAR regions by the original models leads to suboptimal performance under the O-P strategy, as reflected by the drop in mIoU of $7.03\%$ and $2.20\%$ for FC-EF and STANet, respectively.
On the other hand, the TTP algorithm, already exhibiting strong baseline performance, shows highly accurate CAR partitioning in its original model: the predicted CAR distribution closely matches that of the ground truth. Consequently, the O-P strategy provides a notable performance boost, improving mIoU and mFscore by $1.64\%$ and $1.76\%$, respectively. These analyses suggest that the effectiveness of the O-P strategy is highly dependent on the quality of the original model’s CAR predictions: its benefits are only realized when the baseline is sufficiently strong, while a poor baseline renders it less meaningful. Therefore, as a transitional approach, O-P has inherent limitations, whereas the MTKD method, which does not rely on coarse CAR estimation, is more robust and generally preferable.
Appendix G Additional Experiments on the SYSU-CD Dataset
The SYSU-CD dataset is a large-scale, relatively recent, and challenging dataset with a wide CAR range. Therefore, we select it as the second benchmark to evaluate the robustness of our proposed methods. Except for changing the patch size to 256 × 256, all other experimental settings are kept identical to those used for experiments on JL1-CD.
Table 8 presents additional quantitative results under different teacher model configurations, including detection metrics for both the change and no-change classes. As summarized in Table 8, our MTKD approach achieves consistent and notable improvements. After applying MTKD, all models, regardless of whether two or three teacher models are used, exhibit increased mIoU and mFscore values. Specifically, the FC-Siam-Diff model trained with two teacher models achieves the highest improvements in mIoU (+3.18%) and mFscore (+3.76%).
In the last four columns of Table 8, we report detection metrics separately for the change and no-change classes. The absolute values correspond to the original models, while the MTKD framework results are presented as relative improvements. We observe more substantial gains in detecting change areas. For instance, in the two-teacher setup, FC-Siam-Diff’s IoU and Fscore on the change class increase by 6.56% and 7.63%, respectively, whereas performance on the no-change class decreases slightly. These results suggest that MTKD enhances sensitivity toward change regions, consistent with our findings on the JL1-CD dataset.
Visual comparisons are shown in Figure 11. Missed detections remain the most common issue for the original models on this dataset. Both the O-P and MTKD strategies effectively mitigate these issues, with MTKD consistently achieving better results. Nonetheless, limitations remain. For example, while MTKD reduces missed detections in CGNet, some areas remain undetected. Similarly, in TinyCD, although both O-P and MTKD reduce missed detections, a few false positives and holes persist. These observations highlight promising directions for future research to further improve change detection performance.
Table 8: Experimental results on SYSU-CD test set (best and second-best results are in bold and underlined, respectively).

### v4

@@ IV-B MTKD Framework
IV-C Loss Function
where $Y(i)$ denotes the ground truth label of pixel $i$.
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\lambda\mathcal{L}_{\text{KD}},$ (5)
TABLE III: Experimental Results on JL1-CD Test Set
- 57.08 61.90 86.40 61.28 - 63.79 69.54 84.77 69.19
O-P 50.05 53.64 95.28 52.00 O-P 60.93 64.72 91.03 65.46
- 61.30 66.03 86.45 66.34 - 66.76 81.71 74.73 74.02
O-P 60.93 64.72 91.03 65.46 O-P 64.56 78.47 72.13 71.25
- 71.25 78.91 84.53 77.33 - 68.97 74.87 85.06 75.25
O-P 71.06 78.37 84.28 77.21 O-P 71.39 78.60 83.36 77.98
IFN MTKD 72.72 80.28 84.66 78.80 SNUNet MTKD 71.12 78.27 84.96 77.56
- 67.22 74.47 83.71 73.37 - 69.47 75.58 84.46 75.57
O-P 69.41 76.29 84.02 75.77 O-P 68.87 74.74 84.90 74.86
BIT MTKD 68.86 75.49 84.71 74.88
ChangeStar
MTKD 69.87 75.80 85.78 76.03
- 64.85 69.18 88.26 70.19 - 73.51 80.46 86.33 79.70
O-P 64.68 69.05 87.23 70.08 O-P 72.58 79.16 86.33 78.79
ChangeStar
MTKD 65.10 70.26 87.69 70.58
ChangeFormer
(MiT-b0)
MTKD 73.78 80.07 86.90 79.98
- 73.05 79.70 86.95 79.22 - 71.04 78.77 83.05 77.74
O-P 73.45 79.19 87.46 79.41 O-P 72.22 79.93 83.49 78.76
ChangeFormer
(MiT-b1)
MTKD 73.92 80.43 86.89 80.18 TinyCD MTKD 72.55 80.98 83.17 79.26
- 63.64 69.77 83.43 69.39 - 74.85 81.84 86.09 80.98
O-P 69.05 76.53 83.05 75.66 O-P 75.29 81.40 87.06 81.32
HANet MTKD 67.67 74.39 84.38 73.92
Changer
(MiT-b0)
MTKD 75.35 81.76 87.18 81.28
- 75.94 81.99 87.74 81.93 - 68.37 75.15 83.43 74.54
O-P 75.42 81.67 87.13 81.43 O-P 70.76 77.42 83.86 77.01
Changer
(MiT-b1)
MTKD 76.15 82.85 86.98 82.13
Changer
MTKD 69.45 77.26 81.50 75.86
- 62.31 69.23 80.91 67.83 - 64.58 71.16 83.12 70.13
O-P 71.80 79.76 83.15 78.23 O-P 70.19 77.43 83.99 76.16
Changer
MTKD 62.96 69.65 81.76 68.52 LightCDNet MTKD 65.99 72.44 83.86 71.48
- 73.37 80.31 85.33 79.65 - 73.54 79.54 87.89 79.47
O-P 72.95 79.71 85.50 79.12 O-P 73.61 79.17 88.10 79.45
CGNet MTKD 73.82 80.32 86.33 79.91 BAN MTKD 73.95 80.26 87.12 79.92
- 75.05 80.24 89.82 80.76
O-P 76.69 83.48 87.27 82.52
TTP MTKD 76.85 82.99 88.05 82.56
(a) FC-EF
(b) FC-Siam-Conc
(c) FC-Siam-Diff
(d) STANet
(e) IFN
(f) SNUNet
(g) BIT
(h) ChangeStar (FarSeg)
(i) ChangeStar (UPerNet)
(j) ChangeFormer (MiT-b0)
(k) ChangeFormer (MiT-b1)
(l) TinyCD
(m) HANet
(n) Changer (MiT-b0)
(o) Changer (MiT-b1)
(p) Changer (ResNet-18)
(q) Changer (ResNeSt-50)
(r) LightCDNet
(s) CGNet
(t) BAN
(u) TTP
V Experiment
V-A Dataset Description
We first conduct experiments on our JL1-CD dataset. To validate the robustness of the proposed MTKD framework, we further perform experiments on the SYSU-CD dataset [31]. The detailed information for these two datasets is as follows:
1) JL1-CD Dataset: As described in Section III, the JL1-CD dataset consists of 5,000 pairs of high-resolution images, with a resolution of 0.5-0.75 meters and image size of 512 × 512. In the competition, the first 4,000 image pairs are used for training, and the remaining 1,000 image pairs are used for testing.
2) SYSU-CD Dataset: As illustrated in Fig. 9b, the SYSU-CD dataset is also a challenging dataset with a very large CAR range, so we choose this dataset as the second benchmark to validate the robustness of our proposed methods. The SYSU-CD dataset contains 12,000 pairs for training, 4,000 pairs for validation, and 4,000 pairs for testing, with each image having a size of 256 × 256.
V-B Benchmark Methods
To comprehensively verify the validity of our JL1-CD dataset and the effectiveness of the proposed methods, we conduct extensive experiments on existing benchmark algorithms. The selected models, along with their corresponding backbones, parameter sizes, and computational complexities, are summarized in Table II. Based on the backbone architecture, the models are categorized into three groups: Alice Blue for CNN-based models, Light Cyan for Transformer-based models, and Lavender Blue for FM-based models, encompassing almost all mainstream architectures. FLOPs is calculated with an input image of size 512 × 512. As shown in the table, the selected models span a wide range of sizes, from lightweight models with fewer than 1M parameters, such as TinyCD and LightCDNet, to the latest SOTA model TTP with over 360M parameters. This wide range of model sizes allows us to verify the universality of the O-P and MTKD methods across models with different backbones and scales.
V-C Evaluation Metrics
mIoU $=\frac{1}{2}\left(\text{IoU}_{0}+\text{IoU}_{1}\right)$ (6)
$=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FP}+\text{FN}}+\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}}\right),$
mPrec $=\frac{1}{2}\left(\text{Prec}_{0}+\text{Prec}_{1}\right)$ (7)
$=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FN}}+\frac{\text{TP}}{\text{TP}+\text{FP}}\right),$
mRec $=\frac{1}{2}\left(\text{Rec}_{0}+\text{Rec}_{1}\right)$ (8)
$=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FP}}+\frac{\text{TP}}{\text{TP}+\text{FN}}\right),$
mFscore $=\frac{1}{2}\left(\text{Fscore}_{0}+\text{Fscore}_{1}\right)$ (9)
$=\frac{1}{2}\sum_{i=0}^{1}\frac{2\times\text{Prec}_{i}\times\text{Rec}_{i}}{\text{Prec}_{i}+\text{Rec}_{i}},$
V-D Implementation Details
@@
    Training begins with a warm-up phase of 1k iterations (5 epochs for TTP), during which the learning rate (LR) is linearly increased from 1e-6 to the initial LR value, as specified in Table II. Afterward, the LR is linearly decayed to 0 as training progresses (TTP employs a CosineAnnealing decay schedule).
@@
For the training of $\mathcal{M}_{T_{S}}$, due to the varying distributions of the change maps and the different magnitudes of $\mathcal{L}_{\text{CE}}$ across models, a grid search is performed to determine the optimal distillation loss weight $\lambda$ for each model. More configuration details are summarized in Table II.
V-E Experimental Results
V-E1 Quantitative Comparison
Table III summarizes the numerical results of mIoU, mRecall (mRec), mPrecision (mPrec), and mF1-score (mFscore) for all methods on the JL1-CD test set, trained under the original, O-P, and MTKD strategies. As shown in the table, the MTKD framework can improve the performance of all the algorithms to a certain extent. After MTKD optimization, the Changer-MiT-b0 and Changer-MiT-b1 models can outperform the original TTP in terms of mIoU. Additionally, the TTP model, after MTKD optimization, shows improvements in mIoU and mFscore by 1.30% and 1.80%, respectively, setting a new SOTA. However, several algorithms, such as STANet, IFN, and Changer-MiT-b1, perform worse under the O-P strategy than under MTKD, and in some cases, even worse than their original models. This mainly occurs because inaccurate CAR estimation by the original model leads to the selection of an inappropriate teacher model, thereby degrading performance. More qualitative and quantitative analysis of the performance degradation caused by O-P are provided in Section V-F. This finding highlights the potential drawbacks of the O-P strategy in certain scenarios, but it also demonstrates that our MTKD method can enhance student models’ capabilities beyond those of the teacher models trained on partitioned datasets.
TABLE IV: Comparison of Detection Results on Change and No-Change Classes
Method Class IoU Rec Prec Fscore
unchanged +0.10 -0.60 +0.65 +0.06
SNUNet
changed +4.21 +7.38 -0.86 +4.54
unchanged +0.07 +0.12 -0.01 +0.05
ChangeFormer
(MiT-b1)
changed +1.68 +1.35 -0.11 +1.86
unchanged +0.30 +0.29 +0.08 +0.20
TinyCD changed +2.72 +4.13 +0.16 +2.85
unchanged +0.02 -0.03 -0.04 -0.06
CGNet changed +0.88 +0.03 +2.04 +0.59
unchanged +0.12 +0.23 -0.09 +0.07
BAN
changed +0.70 +1.21 -1.46 +0.82
unchanged +0.23 -0.19 +0.45 +0.20
TTP changed +3.36 +5.69 -3.99 +3.39
V-E2 Visual Comparison
In remote sensing CD, four main challenges are typically encountered: false alarms, missed detections, internal consistency, and boundary completeness. To visually assess the effectiveness of the O-P and MTKD strategies in addressing these challenges, we present the visual results of several algorithms in Fig. 4, where red indicates missed detections and blue indicates false alarms. It can be observed that missed detections, particularly for small object changes, occur most frequently across the original algorithms. Certain algorithms (e.g., Changer-ResNet-18, CGNet, and TTP) exhibit false alarms as well. Quantitative analysis reveals that both strategies effectively reduce false alarm and missed detection rates while improving internal consistency and boundary precision.
(a) FC-EF
(b) STANet
(c) TTP
V-E3 Results on Different CAR Partitions
Fig. 5 summarizes the mIoU performance of all benchmark methods across different CAR partitions on the JL1-CD test set. The images are sorted by CAR in ascending order and divided into five equal partitions. Each bar in the figure represents the lower and upper bounds of CAR for each partition, while the line graph indicates the mIoU across the different partitions. The figure indicates that O-P and MTKD yield more notable improvements in partitions with low CAR values, while their impact on high-CAR partitions is generally modest and may even be negative in a few cases (e.g., ChangeFormer-MiT-b1 and TTP). Taking HANet (CNN), ChangeFormer-MiT-b1 (Transformer), and TTP (FM) as representatives, in the first partition of the test set (CAR range 0.00% to 0.88%), O-P improves mIoU by 8.15%, 3.21%, and 6.55%, respectively, while MTKD improves it by 4.42%, 3.36%, and 2.85%. This trend is desirable: original models already perform well on large-scale changes (with mIoU often exceeding 70% in the highest CAR partition), but struggle with small-scale changes (where mIoU frequently falls below 20% in the lowest CAR partition). These results show that by partitioning small CAR regions in the O-P strategy and transferring the knowledge from the corresponding teacher model to the student model in MTKD, our methods achieve substantial improvements in the accuracy of detecting tiny changes.
V-E4 Results on Foreground and Background
@@
We further select some algorithms and compare their detection performance on change and no-change regions. The experimental results are summarized in Table IV. Taking the original models as baselines, we find that MTKD yields consistently larger IoU and Fscore gains on change areas compared to no-change areas, indicating that the MTKD framework is more sensitive to the detection of change areas. Interestingly, we observe that MTKD tends to improve recall more than precision. For instance, the recall of the change region in the MTKD-enhanced TTP model improves by 5.69%, while precision drops by 3.99%. Nevertheless, the final F1-score still increases by 3.39%. This suggests that MTKD effectively reduces the model’s miss rate, and although it may increase false alarms for some algorithms, the overall performance gains are evident. More experimental results on the JL1-CD dataset are provided in Appendix -B.
TABLE V: Impact of Different Numbers of Teacher Models
Method Strategy $\mathcal{M}_{T}$ mIoU mFscore
3 75.29 (+0.44) 81.32 (+0.34)
O-P 2 75.44 (+0.59) 81.51 (+0.53)
3 75.35 (+0.50) 81.28 (+0.30)
Changer
(MiT-b0)
MTKD 2 75.72 (+0.87) 81.66 (+0.68)
3 72.95 (-0.42) 79.12 (-0.53)
O-P 2 73.56 (+0.19) 79.92 (+0.27)
3 73.82 (+0.45) 79.91 (+0.26)
CGNet MTKD 2 73.78 (+0.41) 79.89 (+0.24)
3 76.69 (+1.64) 82.52 (+1.76)
O-P 2 76.65 (+1.60) 82.49 (+1.73)
3 76.85 (+1.80) 82.56 (+1.80)
TTP MTKD 2 76.31 (+1.26) 82.22 (+1.46)
TABLE VI: Experimental Results on SYSU-CD Test Set
Method Strategy mIoU mRec mPrec mFscore
- 70.13 79.20 83.41 77.80
O-P 70.25 79.28 83.88 77.81
SNUNet MTKD 70.61 79.85 83.15 78.29
- 71.19 79.29 87.43 78.41
O-P 71.76 79.84 86.62 78.80
ChangeFormer
(MiT-b0)
MTKD 72.16 80.63 86.63 79.44
- 76.09 84.59 87.08 82.72
O-P 75.97 84.54 86.20 82.66
TTP MTKD 76.11 83.93 87.77 82.77
V-F Limitations of O-P
Whether in Table III or Fig. 5, we observe that some models perform worse with the O-P strategy even compared to their original versions. This can be attributed to partitioning errors caused by wrong CAR estimation, which may lead to suboptimal model selection during the second-stage inference. A deeper analysis of failure scenarios under the O-P strategy is provided. Fig. 6 presents the classification matrix, CAR distribution comparison, and scatter plot between the coarse-predicted CAR of the original models and the GT CAR under the O-P strategy for FC-EF, STANet, and TTP. For FC-EF, the scatter plot reveals a large number of points where high-GT-CAR samples are mispredicted to have almost zero CAR, which indicates a severe issue of missed detections. This observation is further corroborated by the confusion matrix, where a substantial portion of “medium” and “large” regions are incorrectly classified as “small”. As a result, the predicted CAR distribution is heavily skewed, with the number of predicted “small” samples far exceeding that of the ground truth. In contrast, STANet exhibits the opposite trend: its original model tends to generate false alarms: approximately one-quarter of “small” samples are misclassified as “medium”, and more “medium” samples are misclassified as “large” than as “small”. Such misclassification of CAR regions by the original models leads to suboptimal performance under the O-P strategy, as reflected by the drop in mIoU of $7.03\%$ and $2.20\%$ for FC-EF and STANet, respectively.
On the other hand, the TTP algorithm, already exhibiting strong baseline performance, shows highly accurate CAR partitioning in its original model: the predicted CAR distribution closely matches that of the ground truth. Consequently, the O-P strategy provides a notable performance boost, improving mIoU and mFscore by $1.64\%$ and $1.76\%$, respectively.
These analyses suggest that the effectiveness of the O-P strategy is highly dependent on the quality of the original model’s CAR predictions: its benefits are only realized when the baseline is sufficiently strong, while a poor baseline renders it less meaningful. Therefore, as a transitional approach, O-P has inherent limitations, whereas the MTKD method, which does not rely on coarse CAR estimation, is more robust and generally preferable.
V-G Robustness of MTKD
V-G1 Impact of Teacher Configurations in MTKD
In the previous subsection, we have demonstrated that the MTKD framework based on three teacher models can effectively improve the performance of various CD models. However, the 3-teacher configuration requires training five versions of the model for a single CD algorithm: $\mathcal{M}_{O}$, $\mathcal{M}_{T_{L}}$, $\mathcal{M}_{T_{M}}$, $\mathcal{M}_{T_{S}}$, and the student model $\mathcal{M}_{S}$. To alleviate the resource consumption associated with training these models and testing under the O-P strategy, we further explore the performance of the MTKD framework with only two teacher models. We set a threshold of $th=0.10$ to divide the dataset into small and large partitions and train the models $\mathcal{M}_{T_{S}}$ and $\mathcal{M}_{T_{L}}$ on these partitions, respectively. Subsequently, the student model $\mathcal{M}_{S}$ is trained using the MTKD framework. Except for the reduced number of teacher models, all other experimental settings remain unchanged. Using the original models as baselines, Table V compares the metrics of models trained with the O-P strategy and the MTKD framework under different numbers of teacher models. The results indicate that two teacher models can still achieve notable performance improvements over the baseline. For Changer-MiT-b0, the two-teacher model setup yields even greater improvements compared to the three-teacher setup. For CGNet, the O-P strategy with two teachers achieves higher mIoU and mFscore, although the MTKD performance is lower in this case. For TTP, the performance of O-P and MTKD under the three teacher model still remains the strongest. These experiments confirm that the O-P strategy and MTKD framework remain effective with fewer teachers.
V-G2 Performance of MTKD on SYSU-CD
We further evaluate the effectiveness of the MTKD framework on the SYSU-CD dataset. CNN-based SNUNet, Transformer-based ChangeFormer-MiT-b0, and FM-based TTP models are selected. As summarized in Table VI, after MTKD optimization, all the models achieve higher mIoU and mFscore values, with the most notable improvement on ChangeFormer-MiT-b0 (mIoU increased by 0.97% and mFscore increased by 1.03%). However, the O-P strategy does not yield better results on TTP compared to its original version. Overall, MTKD demonstrates substantial effectiveness on the SYSU-CD dataset, further validating its robustness. More numerical results on the SYSU-CD dataset are provided in Appendix -C.
Visual comparisons are shown in Fig. 7. Missed detections remain the most common issue for the original models on this dataset. Both the O-P and MTKD strategies effectively mitigate these issues, with MTKD consistently achieving better results. Nonetheless, limitations remain. For example, while MTKD reduces missed detections in CGNet, some areas remain undetected. Similarly, in TinyCD, although both O-P and MTKD reduce missed detections, a few false positives and holes persist. These observations highlight promising directions for future research to further improve CD performance.
VI Conclusion
In this work, we present JL1-CD, a new benchmark dataset for CD, which substantially augments existing datasets by providing sub-meter resolution, a diverse array of change types, and a large dataset scale. The MTKD framework is further proposed, which consistently improves the performance of mainstream CD algorithms with various architectures and parameter scales, without incurring additional computational cost during inference.
The proposed methods still present certain limitations. Firstly, the O-P approach heavily relies on the performance of the original model and significantly increases resource consumption during inference. MTKD addresses these issues, but a potential limitation of MTKD is that the number of teacher models and the CAR thresholds are determined empirically. Additionally, the resource consumption during the three-stage training process is also considerable.
Future work includes adaptive selection of the number of teacher models and partition thresholds based on different scenarios to further enhance the flexibility and performance of the MTKD framework. Furthermore, optimizing the training process to achieve a balance between accuracy and training efficiency is another avenue for improvement.
(a) JL1-CD
(b) SYSU-CD
Figure 9: CAR distribution of the JL1-CD and SYSU-CD datasets.
-A CAR Distribution
Fig. 8 presents sample images from the JL1-CD dataset with varying values of the Change Area Ratio (CAR). We further provide a quantitative analysis of the CAR distributions for the JL1-CD and SYSU-CD datasets in Fig. 9. The CAR spans from 0 to 100% in these datasets, which poses significant challenges for model learning, and it serves as the primary motivation behind our proposed O-P strategy and MTKD framework.
-B Additional Experiments on the JL1-CD Dataset
We summarize the quantitative results of 21 models on both change and no-change regions in Table VII to provide a comprehensive evaluation of the JL1-CD dataset as a new benchmark and to further validate the effectiveness and generality of the MTKD framework. All models exhibit improvements in IoU and Fscore for change regions after applying MTKD.
@@ VI Conclusion
TABLE VII: Experimental Results on JL1-CD Test Set (to Better Illustrate the Impact of the O-P Strategy and MTKD Framework on the Detection Performance of Change and No-Change Classes, Only the Relative Improvements in Evaluation Metrics Compared to the Original Baseline Model are Reported)
@@
BIT MTKD changed +2.86 +1.37 +2.09 +2.79
ChangeStar
MTKD changed +0.65 +0.51 +2.44 +0.82
@@
ChangeStar
MTKD changed +0.57 +2.31 -1.20 +0.81
ChangeFormer
(MiT-b0)
MTKD changed +0.28 -0.60 +0.69 +0.40
@@
ChangeFormer
(MiT-b1)
MTKD changed +1.68 +1.35 -0.11 +1.86 TinyCD MTKD changed +2.72 +4.13 +0.16 +2.85
@@
HANet MTKD changed +7.27 +8.96 +1.29 +8.53
Changer
(MiT-b0)
MTKD changed +0.80 -0.48 +2.32 +0.41
@@
Changer
(MiT-b1)
MTKD changed +0.41 +1.63 -1.47 +0.42
Changer
(ResNet-18)
MTKD changed +2.07 +4.57 -4.26 +2.57
@@
Changer
(ResNeSt-50)
MTKD changed +0.89 +0.36 +1.77 +1.12 LightCDNet MTKD changed +2.19 +1.70 +1.81 +2.25
@@
TABLE VIII: Experimental Results on SYSU-CD Test Set (Best and Second-Best Results are in Bold and Underlined, Respectively)
@@
-C Additional Experiments on the SYSU-CD Dataset
The SYSU-CD dataset is a large-scale and challenging dataset with a wide CAR range. Therefore, we select it as the second benchmark to evaluate the robustness of our proposed methods. Except for changing the patch size to 256 × 256, all other experimental settings are kept identical to those used for experiments on JL1-CD.
Table VIII presents additional quantitative results under different teacher model configurations, including detection metrics for both the change and no-change classes. As summarized in Table VIII, our MTKD approach achieves consistent and notable improvements. After applying MTKD, all models, regardless of whether two or three teacher models are used, exhibit increased mIoU and mFscore values. Specifically, the FC-Siam-Diff model trained with two teacher models achieves the highest improvements in mIoU (+3.18%) and mFscore (+3.76%).
In the last four columns of Table VIII, we report detection metrics separately for the change and no-change classes. The absolute values correspond to the original models, while the MTKD framework results are presented as relative improvements. We observe more substantial gains in detecting change areas. For instance, in the two-teacher setup, FC-Siam-Diff’s IoU and Fscore on the change class increase by 6.56% and 7.63%, respectively, whereas performance on the no-change class decreases slightly. These results suggest that MTKD enhances sensitivity toward change regions, consistent with our findings on the JL1-CD dataset.